OpenStudy (anonymous):

A line has slope 8 and (3, -1) is on the line. Which is the correct equation that describes this line in slope-intercept form?

OpenStudy (bahrom7893):

y+1=8(x-3) y = 8x -25

OpenStudy (anonymous):

8(x-3)=y+1

OpenStudy (anonymous):

y=8x-25 -8x+y+28=0